# Joint Proceedings of the ACM IUI Workshops 2026, March 23-26, 2026, Paphos, Cyprus

### v1

@@ 1 Introduction
The advent of large language models has triggered a rapid evolution in software development tools, moving from simple code completion to sophisticated, agentic systems. A new frontier in this space is the “prompt-to-app” agent, which aims to generate entire full-stack, deployable web applications from a single natural language description. These systems promise to dramatically lower the barrier to software creation, enabling "citizen developers" and accelerating prototyping for experienced engineers. However, this promise is tempered by a significant lack of rigorous evaluation. How do we know if these generated applications are good? Are they functional? Do they produce results that real users can interact with?
@@
The benchmark we establish provides a critical tool for developers and researchers to navigate the "prompt-to-app" landscape and identifies key areas for future innovation.
@@ 2 Background and Related Work
The evaluation of generative AI has long been a challenge, particularly as outputs become more complex and open-ended. In fields like image and text generation, metrics have evolved from simple pixel-wise or n-gram-based comparisons to more sophisticated perceptual and semantic measures. Recent work has emphasized that evaluating generative AI systems becomes increasingly difficult as outputs grow more open-ended. Studies have shown that traditional pixel-based or token-based similarity metrics fail to capture human-perceived quality or usability [desolda2025hcai, clinical-bertscore, batool2025usability]. This has motivated a shift toward human-centered evaluation frameworks that incorporate real user judgments, interaction behavior, and trust formation.
@@ 3 Benchmark Methodology & Experimental Setup
To assess qualitative aspects of user experience, we employed a two-stage evaluation design using human raters. In the Isolated stage, raters evaluated each application independently using direct five-point scales (1=very poor, 5=very good) on clarity and ease of use. In the Comparison stage, raters viewed both applications side-by-side and provided comparative judgments on a five-point preference scale (1=strongly prefer App A, 3=no preference, 5=strongly prefer App B): visual appeal, visual appropriateness, comparative ease, and trust. At both stages, we recorded whether applications successfully appeared on screen to account for technical failures.
@@ 5.2 Limitations
We evaluated four prompt-to-app systems across a curated set of 96 prompts spanning multiple domains, difficulty levels, specificity, and complexity. Although this prompt set is intentionally diverse, it does not cover the full space of real-world application requirements. As prompt-to-app systems mature, expanding the taxonomy of tasks and incorporating more domain-specific constraints will be essential for broader generalization.
@@ 6 Conclusion and Future Work
The comparative evaluation of agentic "prompt-to-app" systems reveals a landscape of high absolute performance with distinct relative preferences. While the systems are not strictly interchangeable, they all demonstrate high utility, with absolute "Clarity" and "Ease of Use" scores between 3.68 and 3.96 on a 5-point scale.
@@
Our results yield two primary conclusions regarding the current state of agent prompt-to-app systems. First, independent assessments are statistically indistinguishable, suggesting that raters struggle to provide reliable reference-free evaluations in this domain. Consequently, pairwise (side-by-side) comparisons are the only reliable methodology for establishing a meaningful performance benchmark. Second, the current ecosystem does not require a trade-off between key attributes; a single system (Firebase Studio) significantly outperformed competitors across all measured axes, including trust, ease of use, and visual appeal. This indicates that the market is not yet commoditized and certain systems currently dominate across the board, though this may shift as the technology evolves.
@@ A.1.1 Firebase Studio
The automation script executed a multi-phase workflow for each prompt: Initialization: The script launched a Chrome browser instance with persistent session storage to maintain authentication state. Prior to processing, the script verified GitHub API permissions by issuing authenticated requests to ensure the target repository owner (user or organization) had valid credentials for repository creation. Prompt Submission: For each prompt, the script navigated to the Firebase Studio workspace interface, located the prompt input field using CSS selectors, and programmatically entered the prompt text. Prompts were submitted by simulating a keyboard "Enter" event. The script then waited for the platform to generate an initial application blueprint. Prototype Generation: After blueprint generation (typically 25 seconds), the script searched for and clicked the "Prototype this App" button within the Firebase Studio interface. This triggered the platform’s full application generation process. The script monitored for completion indicators for up to 60 attempts at 2-second intervals. Code Export: Once generation completed, the script transitioned the workspace to developer mode, opened an integrated terminal via the VS Code command palette (F1 → "Terminal: Create New Terminal"), and executed a sequence of Git commands to initialize a repository, configure credentials, and push the generated code to a pre-created GitHub repository. Workspace Cleanup: Upon successful export verification (confirmed via GitHub API commit check), the script deleted the Firebase Studio workspace to avoid accumulating workspace clutter and to maintain consistent state for subsequent generations. Concurrency and Rate Limiting. The script processed prompts in batches of 2 concurrent tabs to balance throughput against platform rate limits. A staggered delay was introduced between prompt submissions within each batch. When rate limit responses were detected (identified by parsing page content for "rate limit" or "too many requests" messages), the script paused for a 3-minute cooldown period before retrying. Error Handling and Recovery. The automation included retry logic for transient failures. If the "Prototype this App" button was not found within the expected timeframe—often indicating a "bugged" workspace state—the script attempted workspace recovery by navigating back to the main workspace page, deleting the problematic workspace, and retrying generation with a fresh workspace. Each prompt was allowed up to 2 workspace recovery attempts before being marked as a persistent failure requiring manual investigation. Execution Details. Apps were generated on or around September 5, 2025 using a paid Firebase Studio account. The automation ran in non-headless mode to allow for initial manual authentication. Generated artifacts were pushed to GitHub repositories under a dedicated organization, with repository names derived from sanitized prompt text. A JSON results file was generated at the end of each run, recording timestamps, success/failure status, workspace IDs, and any errors encountered for each prompt.
@@ A.5 Filtering for Quality Responses
    Page 1: Isolated Task for App A. The rater is shown the front page of Wikipedia and given a specific functional task: "Search for information about Albert Einstein."
@@
    Page 1: Isolated Task for App A, expected responses: "Yes" the app appeared, the clarity of the apps purpose should be neutral to positive, and the ease of the task should be neural to positive.
@@
    Page 2: Isolated Task for App B, expected responses: "No" the app did not appear. We did not consider the ratings for the clarity of the app’s purpose or the ease of the task in the smoke test.
@@
    Page 3: Side by Side Preference, expected responses: Report that "Only A" appeared. We did not consider comparative ratings in the smoke test.
@@ A.6 Survey Details
Visual Appropriateness Which app has a more appropriate visual style?
@@ A.11 Prompt Dataset and Taxonomy Examples
This appendix provides a detailed overview of the prompt dataset used to evaluate the agentic "prompt-to-app" systems. To ensure a rigorous and diverse assessment, all prompts were categorized according to a multi-dimensional taxonomy involving Domain, Difficulty, Specificity, and Complexity. This structured approach allows for the evaluation of system performance across varied real-world scenarios, ranging from static informational sites to complex full-stack applications.
@@
    We need a trustworthy and simple application for individuals managing ongoing medication schedules, such as those with chronic conditions or the elderly living independently. The goal is to create a "digital pillbox" that remembers their daily medications. This tool must be incredibly straightforward, requiring no technical skill to operate. The core idea is that a user, perhaps with a family member’s help initially, can set up their list of medications once, and it will always be there when they open the app on their computer or tablet. This persistence is the key feature. The design should prioritize clarity and readability above all else. Think large fonts, clear separation of elements, and obvious controls. While it needs to save data, it should do so locally on the user’s device to maintain privacy and avoid the complexity of accounts and logins. The user should feel a sense of security and control, knowing their personal medication list is stored only where they can access it. The specific layout and features for how they add, view, and manage this list are up to your interpretation.
@@
    You are to build a persistent, client-side "Daily Symptom Diary". This application will allow a user to track their symptoms over time on their own device. The primary goal is to provide users with a simple tool to log how they are feeling each day and save this history for future reference, such as for an upcoming doctor’s appointment. The application must save all entries to the browser’s ‘localStorage‘, ensuring that the user’s diary is preserved between sessions. The interface should be straightforward, allowing users to easily add new entries and view their past logs in a clear, chronological order.
    #### Core Functionality * **New Entry Form**: The main page should feature a simple form to add a new symptom log. * A text input field for "Symptom" (e.g., "Migraine"). * A number input or slider for "Severity" (on a scale of 1 to 10). * A ‘textarea‘ for optional "Notes". * An "Add to Diary" button. * **Symptom History**: * Below the form, a section must display all past diary entries. * Each entry in the history must display the date it was logged, the symptom name, the severity, and the notes. * The entries must be sorted in reverse chronological order (most recent first). * A "Delete" button must be present for each entry, allowing the user to remove it from the diary. * **Data Persistence**: * When the "Add to Diary" button is clicked, the new entry (including an automatic timestamp) is saved to an array in ‘localStorage‘. * When the app loads, it must read the array from ‘localStorage‘ and render the complete symptom history. * Deleting an entry must also update the array in ‘localStorage‘.
@@
    Your task is to create a "Single-Session Pill Reminder," a simple, client-side web application. This tool is designed for users who need a quick, temporary reminder without wanting to create an account or store personal medical information. The core purpose is to set a one-time alert for a medication during their current browser session. All information should be cleared when the page is reloaded, ensuring complete privacy. This is crucial for users on public or shared computers. The application must have a clean, intuitive, and accessible interface that is easy for anyone, including the elderly or less tech-savvy individuals, to use without instructions. Think of it as a digital sticky note for medication.
    #### Core Functionality The application will have a single primary function: setting a one-time reminder. * **Input Form**: The main interface will consist of a form with the following fields: * A text input field labeled "Medication Name" for the user to type the name of the pill (e.g., "Ibuprofen"). * A time input field labeled "Time" for the user to select the time for the reminder. * A button labeled "Set Reminder". * **Reminder Logic**: * When the "Set Reminder" button is clicked, the application should validate that both a name and a future time have been entered. * Upon successful validation, the app should display a confirmation message below the form, such as: "Reminder set for Ibuprofen at 4:30 PM." * At the exact time specified by the user, the browser must trigger a standard ‘alert()‘ dialog box. The alert should read: "Time to take your [Medication Name]!"
@@
    Build a comprehensive "Chronic Condition Management Platform." This is a full-stack web application with a secure backend, designed for patients with chronic illnesses to track their symptoms, treatments, and lifestyle factors over the long term. The platform will require user accounts and will securely store all patient-entered data. Its core purpose is to provide a holistic view of the patient’s health, enabling both the patient and their care team to identify triggers, track treatment efficacy, and make more informed health decisions. The system must support the logging of multiple data types, integrate with external APIs for environmental data, and provide a powerful analytics dashboard.
@@
    Develop an advanced, single-session "Medication Adherence Visualizer". This is a sophisticated clinical tool for healthcare professionals to analyze a patient’s medication schedule during a consultation. The app will not have a database or save any information; its purpose is to perform a complex, one-time analysis and visualization of an imported medication schedule. The professional will import a structured dataset (e.g., a CSV file) representing a patient’s prescribed vs. actual intake times over a period. The application must then process this data and render several interactive charts and graphs to highlight patterns of adherence and non-adherence. The tool is designed for in-depth analysis within a single session, with all data being discarded upon refresh.
    #### Core Functionality * **Data Import**: * An "Import Schedule" button that opens a file selector. * The app must be able to parse a CSV file with the following columns: ‘medication_name‘, ‘prescribed_time‘, ‘actual_intake_time‘. * **Visualization Dashboard**: Once the data is imported, the app must render a dashboard with at least two of the following visualizations: * **Adherence Timeline Chart**: A chart that plots prescribed times vs. actual intake times over the period, showing deviations clearly. * **Time-of-Day Heatmap**: A heatmap showing at which hours of the day adherence is highest and lowest. * **Compliance Score Calculation**: A prominently displayed overall compliance percentage, calculated as the number of on-time doses divided by the total number of prescribed doses. * **Interactivity**: The charts must be interactive. For instance, hovering over a data point on the timeline should show detailed information about that specific dose.

### v2

@@ 1 Introduction
The advent of large language models has triggered a rapid evolution in software development tools, moving from simple code completion to sophisticated, agentic systems. A new frontier in this space is the “prompt-to-app” agent, which aims to generate entire full-stack, deployable web applications from a single natural language description. These systems promise to dramatically lower the barrier to software creation, enabling “citizen developers” and accelerating prototyping for experienced engineers. However, this promise is tempered by a significant lack of rigorous evaluation. How do we know if these generated applications are good? Are they functional? Do they produce results that real users can interact with?
@@
The benchmark we establish provides a critical tool for developers and researchers to navigate the “prompt-to-app” landscape and identifies key areas for future innovation.
@@ 2 Background and Related Work
The evaluation of generative AI has long been a challenge, particularly as outputs become more complex and open-ended. In fields like image and text generation, metrics have evolved from simple pixel-wise or n-gram-based comparisons to more sophisticated perceptual and semantic measures. Recent work has emphasized that evaluating generative AI systems becomes increasingly difficult as outputs grow more open-ended [Anetal2024]. Studies have shown that traditional pixel-based or token-based similarity metrics fail to capture human-perceived quality or usability [desolda2025hcai, clinical-bertscore, batool2025usability]. This has motivated a shift toward human-centered evaluation frameworks that incorporate real user judgments, interaction behavior, and trust formation.
@@ 3 Benchmark Methodology & Experimental Setup
To assess qualitative aspects of user experience, we employed a two-stage evaluation design using human raters. In the Isolated stage, raters evaluated each application independently using direct five-point scales (1=very poor, 5=very good) on clarity and ease of use. In the Comparison stage, raters viewed both applications side-by-side and provided comparative judgments on a five-point preference scale (1=strongly prefer App A, 3=no preference, 5=strongly prefer App B): visual appeal (aesthetic attractiveness), visual appropriateness (whether the design suits the task context), comparative ease, and trust. At both stages, we recorded whether applications successfully appeared on screen to account for technical failures.
@@ 5.2 Limitations
We evaluated three prompt-to-app systems across a curated set of 96 prompts spanning multiple domains, difficulty levels, specificity, and complexity. Although this prompt set is intentionally diverse, it does not cover the full space of real-world application requirements. As prompt-to-app systems mature, expanding the taxonomy of tasks and incorporating more domain-specific constraints will be essential for broader generalization.
@@ 6 Conclusion and Future Work
The comparative evaluation of agentic “prompt-to-app” systems reveals a landscape of high absolute performance with distinct relative preferences. While the systems are not strictly interchangeable, they all demonstrate high utility, with absolute “Clarity” and “Ease of Use” scores between 3.68 and 3.96 on a 5-point scale.
@@
Our results yield two primary conclusions regarding the current state of agent “prompt-to-app” systems. First, independent assessments are statistically indistinguishable, suggesting that raters struggle to provide reliable reference-free evaluations in this domain. Consequently, pairwise (side-by-side) comparisons are the only reliable methodology for establishing a meaningful performance benchmark. Second, the current ecosystem does not require a trade-off between key attributes; a single system (Firebase Studio) significantly outperformed competitors across all measured axes, including trust, ease of use, and visual appeal. This indicates that the market is not yet commoditized and certain systems currently dominate across the board, though this may shift as the technology evolves.
@@ A.1.1 Firebase Studio
The automation script executed a multi-phase workflow for each prompt: Initialization: The script launched a Chrome browser instance with persistent session storage to maintain authentication state. Prior to processing, the script verified GitHub API permissions by issuing authenticated requests to ensure the target repository owner (user or organization) had valid credentials for repository creation. Prompt Submission: For each prompt, the script navigated to the Firebase Studio workspace interface, located the prompt input field using CSS selectors, and programmatically entered the prompt text. Prompts were submitted by simulating a keyboard “Enter” event. The script then waited for the platform to generate an initial application blueprint. Prototype Generation: After blueprint generation (typically 25 seconds), the script searched for and clicked the “Prototype this App” button within the Firebase Studio interface. This triggered the platform’s full application generation process. The script monitored for completion indicators for up to 60 attempts at 2-second intervals. Code Export: Once generation completed, the script transitioned the workspace to developer mode, opened an integrated terminal via the VS Code command palette (F1 → “Terminal: Create New Terminal”), and executed a sequence of Git commands to initialize a repository, configure credentials, and push the generated code to a pre-created GitHub repository. Workspace Cleanup: Upon successful export verification (confirmed via GitHub API commit check), the script deleted the Firebase Studio workspace to avoid accumulating workspace clutter and to maintain consistent state for subsequent generations. Concurrency and Rate Limiting. The script processed prompts in batches of 2 concurrent tabs to balance throughput against platform rate limits. A staggered delay was introduced between prompt submissions within each batch. When rate limit responses were detected (identified by parsing page content for “rate limit” or “too many requests” messages), the script paused for a 3-minute cooldown period before retrying. Error Handling and Recovery. The automation included retry logic for transient failures. If the “Prototype this App” button was not found within the expected timeframe—often indicating a “bugged” workspace state—the script attempted workspace recovery by navigating back to the main workspace page, deleting the problematic workspace, and retrying generation with a fresh workspace. Each prompt was allowed up to 2 workspace recovery attempts before being marked as a persistent failure requiring manual investigation. Execution Details. Apps were generated on or around September 5, 2025 using a paid Firebase Studio account. The automation ran in non-headless mode to allow for initial manual authentication. Generated artifacts were pushed to GitHub repositories under a dedicated organization, with repository names derived from sanitized prompt text. A JSON results file was generated at the end of each run, recording timestamps, success/failure status, workspace IDs, and any errors encountered for each prompt.
@@ A.5 Filtering for Quality Responses
    Page 1: Isolated Task for App A. The rater is shown the front page of Wikipedia and given a specific functional task: “Search for information about Albert Einstein.”
@@
    Page 1: Isolated Task for App A, expected responses: “Yes” the app appeared, the clarity of the apps purpose should be neutral to positive, and the ease of the task should be neural to positive.
@@
    Page 2: Isolated Task for App B, expected responses: “No” the app did not appear. We did not consider the ratings for the clarity of the app’s purpose or the ease of the task in the smoke test.
@@
    Page 3: Side by Side Preference, expected responses: Report that “Only A” appeared. We did not consider comparative ratings in the smoke test.
@@ A.6 Survey Details
Visual Appropriateness Which app has a more appropriate visual style for the task?
@@ A.11 Prompt Dataset and Taxonomy Examples
This appendix provides a detailed overview of the prompt dataset used to evaluate the agentic “prompt-to-app” systems. To ensure a rigorous and diverse assessment, all prompts were categorized according to a multi-dimensional taxonomy involving Domain, Difficulty, Specificity, and Complexity. This structured approach allows for the evaluation of system performance across varied real-world scenarios, ranging from static informational sites to complex full-stack applications.
@@
    We need a trustworthy and simple application for individuals managing ongoing medication schedules, such as those with chronic conditions or the elderly living independently. The goal is to create a “digital pillbox” that remembers their daily medications. This tool must be incredibly straightforward, requiring no technical skill to operate. The core idea is that a user, perhaps with a family member’s help initially, can set up their list of medications once, and it will always be there when they open the app on their computer or tablet. This persistence is the key feature. The design should prioritize clarity and readability above all else. Think large fonts, clear separation of elements, and obvious controls. While it needs to save data, it should do so locally on the user’s device to maintain privacy and avoid the complexity of accounts and logins. The user should feel a sense of security and control, knowing their personal medication list is stored only where they can access it. The specific layout and features for how they add, view, and manage this list are up to your interpretation.
@@
    You are to build a persistent, client-side “Daily Symptom Diary”. This application will allow a user to track their symptoms over time on their own device. The primary goal is to provide users with a simple tool to log how they are feeling each day and save this history for future reference, such as for an upcoming doctor’s appointment. The application must save all entries to the browser’s ‘localStorage‘, ensuring that the user’s diary is preserved between sessions. The interface should be straightforward, allowing users to easily add new entries and view their past logs in a clear, chronological order.
    #### Core Functionality * **New Entry Form**: The main page should feature a simple form to add a new symptom log. * A text input field for “Symptom” (e.g., “Migraine”). * A number input or slider for “Severity” (on a scale of 1 to 10). * A ‘textarea‘ for optional “Notes”. * An “Add to Diary” button. * **Symptom History**: * Below the form, a section must display all past diary entries. * Each entry in the history must display the date it was logged, the symptom name, the severity, and the notes. * The entries must be sorted in reverse chronological order (most recent first). * A “Delete” button must be present for each entry, allowing the user to remove it from the diary. * **Data Persistence**: * When the “Add to Diary” button is clicked, the new entry (including an automatic timestamp) is saved to an array in ‘localStorage‘. * When the app loads, it must read the array from ‘localStorage‘ and render the complete symptom history. * Deleting an entry must also update the array in ‘localStorage‘.
@@
    Your task is to create a “Single-Session Pill Reminder”, a simple, client-side web application. This tool is designed for users who need a quick, temporary reminder without wanting to create an account or store personal medical information. The core purpose is to set a one-time alert for a medication during their current browser session. All information should be cleared when the page is reloaded, ensuring complete privacy. This is crucial for users on public or shared computers. The application must have a clean, intuitive, and accessible interface that is easy for anyone, including the elderly or less tech-savvy individuals, to use without instructions. Think of it as a digital sticky note for medication.
    #### Core Functionality The application will have a single primary function: setting a one-time reminder. * **Input Form**: The main interface will consist of a form with the following fields: * A text input field labeled “Medication Name” for the user to type the name of the pill (e.g., “Ibuprofen”). * A time input field labeled “Time” for the user to select the time for the reminder. * A button labeled “Set Reminder”. * **Reminder Logic**: * When the “Set Reminder” button is clicked, the application should validate that both a name and a future time have been entered. * Upon successful validation, the app should display a confirmation message below the form, such as: “Reminder set for Ibuprofen at 4:30 PM.” * At the exact time specified by the user, the browser must trigger a standard ‘alert()‘ dialog box. The alert should read: “Time to take your [Medication Name]!”
@@
    Build a comprehensive “Chronic Condition Management Platform.” This is a full-stack web application with a secure backend, designed for patients with chronic illnesses to track their symptoms, treatments, and lifestyle factors over the long term. The platform will require user accounts and will securely store all patient-entered data. Its core purpose is to provide a holistic view of the patient’s health, enabling both the patient and their care team to identify triggers, track treatment efficacy, and make more informed health decisions. The system must support the logging of multiple data types, integrate with external APIs for environmental data, and provide a powerful analytics dashboard.
@@
    Develop an advanced, single-session “Medication Adherence Visualizer”. This is a sophisticated clinical tool for healthcare professionals to analyze a patient’s medication schedule during a consultation. The app will not have a database or save any information; its purpose is to perform a complex, one-time analysis and visualization of an imported medication schedule. The professional will import a structured dataset (e.g., a CSV file) representing a patient’s prescribed vs. actual intake times over a period. The application must then process this data and render several interactive charts and graphs to highlight patterns of adherence and non-adherence. The tool is designed for in-depth analysis within a single session, with all data being discarded upon refresh.
    #### Core Functionality * **Data Import**: * An “Import Schedule” button that opens a file selector. * The app must be able to parse a CSV file with the following columns: ‘medication_name‘, ‘prescribed_time‘, ‘actual_intake_time‘. * **Visualization Dashboard**: Once the data is imported, the app must render a dashboard with at least two of the following visualizations: * **Adherence Timeline Chart**: A chart that plots prescribed times vs. actual intake times over the period, showing deviations clearly. * **Time-of-Day Heatmap**: A heatmap showing at which hours of the day adherence is highest and lowest. * **Compliance Score Calculation**: A prominently displayed overall compliance percentage, calculated as the number of on-time doses divided by the total number of prescribed doses. * **Interactivity**: The charts must be interactive. For instance, hovering over a data point on the timeline should show detailed information about that specific dose.